# LLMR: Real-time Prompting of Interactive Worlds using Large Language Models

### 19 Dec 2023 v2

@@ Abstract.
Figure 1. Examples of diverse use cases and functionalities enabled by Large Language Model for Mixed Reality (LLMR). A: Creation of a kitchen scene from scratch using Unity primitives. B: Prompting and drawing objects into existence via multi-modal interactions. C: Integration with external plugins like SketchFab to create high-fidelity scenes and prompting skills like animation creation. D: Edit existing VR scenes. E: Automated generation of instructional guides and QA with scene knowledge. F: Cross-platform compatibility and external sensor integration.
@@ 1. Introduction
LLMR is an orchestration of an ensemble of specialized GPTs. At its center is the Builder GPT serving as an architect of C# Unity code for crafting interactive scenes. However, the multitude of tasks falling under virtual world creation renders a standalone coder insufficient. For instance, the ability to meaningfully modify an existing virtual world necessitates a profound semantic understanding of the scene. As humans, we have the ability to infer the properties of objects in the world and can refer to objects in the environment using demonstratives. To simulate the benefits of perceptual access, we incorporated the Scene Analyzer GPT. It generates a comprehensive summary of scene objects, offering detailed information when requested, including aspects like size, color, and the functionalities of interactive tools previously generated by LLMR. We also implemented the Skill Library GPT that determines the relevant skills that are needed for the Builder to accomplish the user’s request. In addition, we have observed that the code generated by the Builder lacks robustness and frequently contains bugs. To remedy this, we introduce the Inspector GPT, which evaluates the Builder’s code against a predefined set of rules. This evaluation acts as a protective measure against compilation and run-time errors before the code is executed via the Compiler.
To illustrate the efficacy of our framework in the creation and editing of virtual scenes, we tested LLMR on two sets of 150 prompts encompassing a wide array of creation and modification tasks. Our findings demonstrate LLMR’s superior performance in contrast to general-purpose LLMs while emphasizing the performance gain achieved with the addition of each module in our pipeline. In particular, LLMR exhibits over 3x reduction in code errors in both an empty and an existing scene, when compared to off-the-shelf GPT-4 (OpenAI, 2023). In the meantime, LLMR can successfully complete sequences of tasks with varying complexities, while keeping the completion time around a minute. These outcomes underscore LLMR’s capacity to execute user instructions in real-time with a higher degree of robustness.
@@
Our paper is organized as follows: we begin by describing prior work and approaches to generating 3D objects and environments for mixed reality in Section 2. In Section 3, we first provide an overview of LLMR followed by details of the function of each module of our framework. We then discuss important extensions of our framework, such as incorporating plugins, memory management, and cross-platform compatibility, in Sections 4,5,6, respectively. We then present a series of exemplar applications in Section 7 to illustrate the wide range of creations enabled by LLMR. Section 8 Numerical Study includes a comprehensive evaluation of our framework against our design goals: high completion rate, real-time execution, robust against complex tasks, and iterative fine-tune ability. We follow the Numerical Study with Section 9 User study that evaluates the quality of LLMR’s output and presents usability feedback. Finally, in Section 10, we discuss the limitations and future work for others to build upon.
@@ 2.1. Generative 3D Assets
The generation of 3D assets has been a significant focus in recent research. The work of Li et al. with 3DDesigner (Li et al., 2022), Jun and Nichol with Shap·E (Jun and Nichol, 2023), and Poole et al. with DreamFusion (Poole et al., 2022) have demonstrated the potential of text guidance and generative models in creating complex and diverse 3D objects. Lin et al. introduce Magic3D (Lin et al., 2023), a high-resolution text-to-3D content creation framework that addresses the limitations of slow optimization and low-resolution output inherent in existing methods like DreamFusion. Recently, Holodiffusion by Karnewar et al. (Karnewar et al., 2023) furthered the conversation by employing diffusion models for 3D generative modeling. The Instruct-NeRF2NeRF method (Haque et al., 2023) and advancements like Pointclip v2 (Zhu et al., 2023) as well as the work of Roberts et al. (Roberts et al., 2022a) have explored the power of prompting techniques in 3D open-world learning. Our approach extends beyond visual appearances to incorporate interactive and behavioral elements into the generated content.
2.2. Interactive 3D Environments
Recently, the creation of interactive 3D environments has been further explored, with contributions from Wang et al. with Voyager (Wang et al., 2023d), Singer et al. with MAV3D (Singer et al., 2023), and Höllein, Lukas, et al. with Text2Room (Höllein et al., 2023). Volum et al. has shown that LLMs can be used to guide NPC interactions with a virtual environment (Volum et al., 2022). Wang et al. also introduced Chat-3D (Wang et al., 2023c), a system that focuses on universal dialogues for 3D scenes, which is further augmented by the work of Hong et al. with 3D-LLM (Hong et al., 2023). New approaches like Oasis (Sra et al., 2017a) and Procedurally Generated Virtual Reality (Sra et al., 2016) add novel perspectives. Thoravi Kumaravel et al. (Thoravi Kumaravel et al., 2019) complement these efforts by focusing on bi-directional mixed-reality telepresence. While these contributions are significant in building interactive 3D spaces, the interplay between AI and mixed reality in these environments remains an open question. Our work tackles this gap by bringing the capabilities of LLMs to a real-time Unity editor for Mixed Reality applications.
@@ 2.3. Editor Support for Mixed Reality Development
Mixed Reality (XR) development has been explored by Hirzle et al. (Hirzle et al., 2023) and Fidalgo et al. (Fidalgo et al., 2023), who provide comprehensive reviews at the intersection of AI and XR. Lindlbauer et al. (Lindlbauer et al., 2019) and Cheng et al. (Cheng et al., 2021) focus on the automatic adaptation of MR interfaces, a line of work that is relevant for multi-user XR experiences, as shown by Mandi et al. with RoCo (Mandi et al., 2023).
2.4. Procedural Generation and Object Compositing
Research by Freiknecht and Effelsberg (Freiknecht and Effelsberg, 2017), Cao et al. (Cao and Johnson, 2023), and Song et al. (Song et al., 2023) has focused on the balance between realism and algorithmic performance. Recent advancements such as Interactive Example-Based Terrain Authoring with Conditional Generative Adversarial Networks by Guérin et al. (Guerin et al., 2017) add a layer of complexity in how terrains can be generated from simple user inputs. DeepSpace introduced a novel method of mood-based texture generation from music (Sra et al., 2017b), adding another layer of complexity to asset generation. Daffara et al. (Daffara et al., 2020) and Rana et al. (Rana et al., 2023) further extended these concepts to include demonstrations and task planning.
2.5. Interpreting Non-Linguistic Information
Lastly, the work of Zhang et al. with MotionGPT (Zhang et al., 2023a), Wu et al.’s work on Embodied Task Planning (Wu et al., 2023) as well as Richardson et al. with TEXTure (Richardson et al., 2023), and Driess et al. with PaLM-E (Driess et al., 2023) has shown the potential of LLMs in generating human motion, texturing 3D shapes, and incorporating real-world sensor modalities, respectively. These efforts are complemented by Xu et al. with XAIR (Xu et al., 2023), which focuses on explainable AI in augmented reality. A comprehensive review of Neural Radiance Field (NeRF) models by Gao et al. (Gao et al., 2022) adds to our understanding of this rapidly growing field and aligns with our approach of enabling LLMs to interpret non-linguistic or non-symbolic information.
@@ 3. LLMR: a framework for generating real-time, interactive 3D worlds using large language models
Figure 2. Large Language Model for Mixed Reality (LLMR) architecture for real-time interactive 3D scene generation. Starting from the left, a user prompt and the existing 3D scene ($\Omega$) are fed into the Planner and Scene Analyzer (SA) modules, respectively. The Planner decomposes the user prompt into a sequence of sub-prompts, while the SA summarizes the current scene elements. These are then integrated with a Skill Library (SL) to guide the Builder module, which generates the appropriate code. The Inspector module iteratively checks the generated code for compilation and run-time errors. Upon receiving the green light from the Inspector, the code is compiled using the Roslyn Compiler and executed in the Unity Engine to produce the desired 3D scene and functionalities as specified by the user.
@@
Large language models are capable code generators, and their ability to synthesize programs have been extensively tested (Austin et al., 2021; Nijkamp et al., 2022; Christopoulou et al., 2022; Gur et al., 2023; Jain et al., 2022). Scripting in a game engine, however, is especially challenging given the multitude of tasks and the complexity of the development environment. For a non-comprehensive list, generating a realistic 3D world may involve object creation, texturing, behavior programming, event scripting, animations, particle effects, lighting, and user interface (Bethke, 2003). Prompting these elements in real time requires a framework that understands the virtual scene, interprets user intention, and generates high-quality code. To this end, we present Large Language Model for Mixed Reality (LLMR), a framework that enables real-time creation and modification of interactive 3D scenes using natural language.
@@
LLMR is an orchestration of language models, each contextualized with a distinct metaprompt to outline its role, as illustrated in Figure 2. A metaprompt is a specially crafted input sequence or context that guides an LLM’s behavior or output, enabling more focused or nuanced responses than standard prompts. We start with the Planner, which breaks down the user’s request into a sequence of appropriately scoped instructions. These instructions, along with a concise summary of the existing scene and extra knowledge for specialized skills, are used as inputs to the central module called Builder, which generates code to fulfill these instructions. In addition, we use a separate Inspector module to check the Builder’s generated code against potential compilation and run-time errors before finally executing the code. Below, we detail each module and explain the design choices that enable various aspects of prompting a virtual world into existence.
@@
The task of generating interactive 3D scenes boils down to generating and executing appropriate code snippets to accomplish the user’s prompt. Formally, denote the user’s request by $u$ and the current 3D world by $\Omega$ (which may be empty), we wish to draw sample $x\sim\mathcal{P}(x|u,\Omega)$, where $\mathcal{P}$ is the distribution of syntactically valid, request-fulfilling code. We then compile and execute $x$ at run-time under the Unity Engine (Unity Technologies, 2005), a development platform for creating virtual scenes that suits our needs.
@@
1:$u$: user’s request; $\Omega:$ current scene; $\mathrm{A}(s|u,\Omega)$: Scene Analyzer; $\mathrm{P}(u_{1},...,u_{N}|u,s)$: Planner; $\mathrm{L}(h|u)$: Skill Library; $\mathrm{B}(x|u,s,h):$ Builder; $\mathrm{I}(r,v|u,s,x)$: Inspector.
@@
[…]
4:$\Omega_{1}\leftarrow\Omega$
5:for $i=1:N$ do
6:     $s_{i}\sim\mathrm{A}(\cdot|u_{i},\Omega_{i})$triangleright Analyze the current scene.
7:     $h_{i}\sim\mathrm{L}(\cdot|u_{i})$triangleright Retrieve required skills, if any.
8:     $x_{i}\leftarrow\textrm{GenerateCodeWithInspection}(u_{i},s_{i},h_{i})$triangleright Generate verified code per the instruction.
9:     $\Omega_{i+1}\leftarrow\textrm{CompileAndRun}(x_{i},\Omega_{i})$triangleright Execute the code to change the world state.
10:end for

### 22 Mar 2024 v3

@@ Abstract.
large language model, mixed reality, spatial reasoning, artificial intelligence
journalyear: 2024copyright: rightsretainedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642579isbn: 979-8-4007-0330-0/24/05ccs: Computing methodologies Spatial and physical reasoningccs: Computing methodologies Multi-agent systems
Figure 1. Examples of diverse use cases and functionalities enabled by the Large Language Model for Mixed Reality (LLMR) framework. A: Creation of a detailed kitchen scene from scratch using Unity primitives. B: Prompting and drawing objects into existence via multi-modal interactions. C: Integration with external plugins like loading objects from Sketchfab to create high-fidelity scenes and special skills like generating animations. D: Prompting edits of existing VR scenes like changing the color of the objects. E: Automated generation of instructional guides and Questioning and Answering about the scene. F: The framework is compatible across platforms and supports the integration of external sensor data.
Examples of diverse use cases and functionalities enabled by Large Language Model for Mixed Reality. This composite image displays six panels, each representing different functionalities of a Mixed Reality framework. A: World Creation - A virtual 3D kitchen constructed with basic geometric shapes such as cubes and rectangles, representing furniture and appliances. B: Multi-modal Interaction - A virtual reality controller with a blue line-rendering trailing from its tip, showing dynamic drawing in a 3D space. C: External Plugin Integration - A virtual room with a wall clock, a picture frame, and furniture, illustrating the incorporation of detailed objects from an external source. D: Edit Existing Worlds - A kitchen scene with a high level of detail, including cabinets, appliances, and decor, demonstrating the ability to modify pre-existing virtual spaces. E: Question and Answer with Scene - A partially disassembled virtual espresso machine with a controller pointing at it, indicating an interactive environment where users can query and receive information. F: Cross-platform and Sensor Integration - A person’s wrist wearing a sensor device, with a virtual whale in front of them, suggesting synchronization between physical movements or biometrics and virtual reality actions. Each panel is labeled with a letter from A to F and includes a brief description of the feature or task demonstrated, such as creating environments from scratch, integrating complex objects, or editing existing virtual spaces for accessibility.
Figure 1. Examples of diverse use cases and functionalities enabled by the Large Language Model for Mixed Reality (LLMR) framework. A: Creation of a detailed kitchen scene from scratch using Unity primitives. B: Prompting and drawing objects into existence via multi-modal interactions. C: Integration with external plugins like loading objects from Sketchfab to create high-fidelity scenes and special skills like generating animations. D: Prompting edits of existing VR scenes like changing the color of the objects. E: Automated generation of instructional guides and Questioning and Answering about the scene. F: The framework is compatible across platforms and supports the integration of external sensor data.
@@ 1. Introduction
LLMR is an orchestration of an ensemble of specialized GPTs. At its center is the Builder GPT serving as an architect of C# Unity code for crafting interactive scenes. However, the multitude of tasks falling under virtual world creation renders a standalone coder insufficient. For instance, the ability to meaningfully modify an existing virtual world necessitates a profound semantic understanding of the scene. As humans, we have the ability to infer the properties of objects in the world and can refer to objects in the environment using demonstratives. To simulate the benefits of perceptual access, we incorporated the Scene Analyzer GPT. It generates a comprehensive summary of scene objects, offering detailed information when requested, including aspects like size, color, and the functionalities of interactive tools previously generated by LLMR. We also implemented the Skill Library GPT that determines the relevant skills that are needed for the Builder to accomplish the user’s request. In addition, we have observed that the code generated by the Builder lacks robustness and frequently contains bugs. To remedy this, we introduce the Inspector GPT, which evaluates the Builder’s code against a predefined set of rules. This evaluation acts as a protective measure against compilation and run-time errors before the code is executed via the Compiler in the Unity Game Engine.
To illustrate the efficacy of our framework in the creation and editing of virtual scenes, we tested LLMR on two sets of 150 prompts encompassing a wide array of creation and modification tasks. Our findings demonstrate LLMR’s superior performance in contrast to general-purpose LLMs while emphasizing the performance gain achieved with the addition of each module in our pipeline. In particular, LLMR exhibits 4x reduction in code errors in both an empty and an existing scene, when compared to off-the-shelf GPT-4 (OpenAI, 2023). In the meantime, LLMR can successfully complete sequences of tasks with varying complexities, while keeping the completion time around a minute. These outcomes underscore LLMR’s capacity to execute user instructions in real time with a higher degree of robustness.
@@
Our paper is organized as follows: we begin by describing prior work and approaches to generating 3D objects and environments for mixed reality in Section 2. In Section 3, we first provide an overview of LLMR followed by details of the function of each module of our framework. We then discuss important extensions of our framework, such as incorporating plugins, memory management, and cross-platform compatibility, in Sections 4,5,6, respectively. We then present a series of exemplar applications in Section 7 to illustrate the wide range of creations enabled by LLMR. Section 8 Numerical Study includes a comprehensive evaluation of our framework against our design goals: high completion rate, real-time execution, robust against complex tasks, and iterative fine-tuning ability. We follow the Numerical Study with Section 9 User study that evaluates the quality of LLMR’s output and presents usability feedback. Finally, in Section 10, we discuss the limitations and future work for others to build upon.
@@ 2.1. Generative 3D Assets
The generation of 3D assets has been a significant focus in recent research. The work of Li et al. with 3DDesigner (Li et al., 2022), Jun and Nichol with Shap·E (Jun and Nichol, 2023), and Poole et al. with DreamFusion (Poole et al., 2022) have demonstrated the potential of text guidance and generative models in creating complex and diverse 3D objects. Lin et al. introduce Magic3D (Lin et al., 2023), a high-resolution text-to-3D content creation framework that addresses the limitations of slow optimization and low-resolution output inherent in existing methods like DreamFusion. Recently, Holodiffusion by Karnewar et al. (Karnewar et al., 2023) furthered the conversation by employing diffusion models for 3D generative modeling. The Instruct-NeRF2NeRF method (Haque et al., 2023) and advancements like Pointclip v2 (Zhu et al., 2023) as well as the work of Roberts et al. (Roberts et al., 2022a) have explored the power of prompting techniques in 3D open-world learning. A comprehensive review of Neural Radiance Field (NeRF) models by Gao et al. (Gao et al., 2022) adds to our understanding of this rapidly growing field and aligns with our approach of enabling LLMs to interpret non-linguistic or non-symbolic information. Our approach extends beyond visual appearances to incorporate interactive and behavioral elements into the generated content.
2.2. Generative Interactive 3D Environments
In addition to generating objects, the creation of interactive 3D environments has been further explored, with contributions from Wang et al. with Voyager (Wang et al., 2023d), Singer et al. with MAV3D (Singer et al., 2023), and Höllein, Lukas, et al. with Text2Room (Höllein et al., 2023). Volum et al. has shown that LLMs can be used to guide NPC interactions with a virtual environment (Volum et al., 2022). Wang et al. also introduced Chat-3D (Wang et al., 2023c), a system that focuses on universal dialogues for 3D scenes, which is further augmented by the work of Hong et al. with 3D-LLM (Hong et al., 2023). New approaches like Oasis (Sra et al., 2017a) and Procedurally Generated Virtual Reality (Sra et al., 2016) add novel perspectives. Recent advancements such as Interactive Example-Based Terrain Authoring with Conditional Generative Adversarial Networks by Guérin et al. (Guerin et al., 2017) add a layer of complexity to how terrains can be generated from simple user inputs. Research by Freiknecht and Effelsberg (Freiknecht and Effelsberg, 2017), Cao et al. (Cao and Johnson, 2023), and Song et al. (Song et al., 2023) has focused on the balance between realism and algorithmic performance. DeepSpace introduced a novel method of mood-based texture generation from music (Sra et al., 2017b), adding another layer of complexity to asset generation. While these contributions are significant in building interactive 3D spaces, the interplay between AI and mixed reality in these environments remains an open question. Our work tackles this gap by bringing the capabilities of LLMs to a real-time Unity editor for Mixed Reality applications.
@@ 2.3. Editor Support for Mixed Reality Development
Mixed Reality (XR) development has been explored by Hirzle et al. (Hirzle et al., 2023) and Fidalgo et al. (Fidalgo et al., 2023), who provide comprehensive reviews at the intersection of AI and XR. Lindlbauer et al. (Lindlbauer et al., 2019) and Cheng et al. (Cheng et al., 2021) focus on the automatic adaptation of MR interfaces, a line of work that is relevant for multi-user XR experiences, as shown by Mandi et al. with RoCo (Mandi et al., 2023). Thoravi Kumaravel et al. (Thoravi Kumaravel et al., 2019) complement these efforts by focusing on bi-directional mixed-reality telepresence. Compared with prior work, we allow users to directly authorize the environment using natural language.
2.4. LLMs Interpreting Spatial, Non-Linguistic Information
Lastly, many have pushed the boundary of LLMs by inputting non-linguistic information (which was not in the training set), such as for visual programming (Zhou et al., 2023) or processing sensor data (Liu et al., 2023b). More related to our work is using LLM to interface with spatial, embodied data. Work of Zhang et al. with MotionGPT (Zhang et al., 2023a), Wu et al.’s work on Embodied Task Planning (Wu et al., 2023) as well as Richardson et al. with TEXTure (Richardson et al., 2023). Daffara et al. (Daffara et al., 2020) and Rana et al. (Rana et al., 2023) further extended these concepts to include demonstrations and task planning. Driess et al. with PaLM-E (Driess et al., 2023) has shown the potential of LLMs in generating human motion, texturing 3D shapes, and incorporating real-world sensor modalities, respectively. These efforts are complemented by Xu et al. with XAIR (Xu et al., 2023), which focuses on explainable AI in augmented reality. We hope LLMR contribute to the improvement of LLM’s capability of spatial reasoning and world understanding.
Figure 2. Large Language Model for Mixed Reality (LLMR) architecture for real-time interactive 3D scene generation. Starting from the left, a user prompt and the existing 3D scene ($\Omega$) are fed into the Planner (P) and Scene Analyzer (SA) modules, respectively. The Planner decomposes the user prompt into a sequence of sub-prompts, while the SA summarizes the current scene elements. These are then integrated with a Skill Library (SL) to guide the Builder (B) module, which generates the appropriate code. The Inspector (I) module iteratively checks the generated code for compilation and run-time errors. Upon receiving the green light from the Inspector, the code is compiled using the Roslyn Compiler and executed in the Unity Engine to produce the desired 3D scene and functionalities as specified by the user.
[…]
Figure 2. Large Language Model for Mixed Reality (LLMR) architecture for real-time interactive 3D scene generation. Starting from the left, a user prompt and the existing 3D scene ($\Omega$) are fed into the Planner (P) and Scene Analyzer (SA) modules, respectively. The Planner decomposes the user prompt into a sequence of sub-prompts, while the SA summarizes the current scene elements. These are then integrated with a Skill Library (SL) to guide the Builder (B) module, which generates the appropriate code. The Inspector (I) module iteratively checks the generated code for compilation and run-time errors. Upon receiving the green light from the Inspector, the code is compiled using the Roslyn Compiler and executed in the Unity Engine to produce the desired 3D scene and functionalities as specified by the user.
@@ 3. LLMR: a framework for generating real-time, interactive 3D worlds using large language models
Large language models are capable code generators, and their ability to synthesize programs has been extensively tested (Austin et al., 2021; Nijkamp et al., 2022; Christopoulou et al., 2022; Gur et al., 2023; Jain et al., 2022). Scripting in a game engine, however, is especially challenging given the multitude of tasks and the complexity of the development environment. For a non-comprehensive list, generating a realistic 3D world may involve object creation, texturing, behavior programming, event scripting, animations, particle effects, lighting, and user interface (Bethke, 2003). Prompting these elements in real time requires a framework that understands the virtual scene, interprets user intention, and generates high-quality code. To this end, we present Large Language Model for Mixed Reality (LLMR), a framework that enables real-time creation and modification of interactive 3D scenes using natural language.
@@
LLMR is an orchestration of language models, each contextualized with a distinct metaprompt to outline its role, as illustrated in Figure 2 and Algorithm 1. A metaprompt is a specially crafted input sequence or context that guides an LLM’s behavior or output, enabling more focused or nuanced responses than standard prompts. We start with the Planner, which breaks down the user’s request into a sequence of appropriately scoped instructions. These instructions, along with a concise summary of the existing scene from the Scene Analyzer and extra knowledge for specialized skills from the Skill Library, are used as inputs to the central module called Builder, which generates code to fulfill these instructions. In addition, we use a separate Inspector module to check the Builder’s generated code against potential compilation and run-time errors before finally executing the code.
@@
The task of generating interactive 3D scenes boils down to generating and executing appropriate code snippets to accomplish the user’s prompt. Formally, denote the user’s request by $u$ and the current 3D world by $\Omega$ (which may be empty), we wish to draw sample $x\sim\mathcal{P}(x|u,\Omega)$, where $\mathcal{P}$ is the distribution of syntactically valid, request-fulfilling code. We then compile and execute $x$ at run-time under the Unity Engine (Unity Technologies, 2005), a development platform for creating virtual scenes that suits our needs. Below, we detail each module and explain the design choices that enable various aspects of prompting a virtual world into existence.
Input : $u$: user’s request; $\Omega$: current scene
Require : $\mathrm{A}(s|u,\Omega)$: Scene Analyzer;
$\mathrm{P}(u_{1},...,u_{N}|u,s)$: Planner;
$\mathrm{L}(h|u)$: Skill Library;
$\mathrm{B}(x|u,s,h)$: Builder;
$\mathrm{I}(r,v|u,s,x)$: Inspector.
$s\sim\mathrm{A}(\cdot|u,\Omega)$;
$(u_{1},...,u_{N})\sim\mathrm{P}(\cdot|u,s)$ /* Decomposes the request into suitable instructions. */
 $\Omega_{1}\leftarrow\Omega$;
for $i=1:N$ do
       $s_{i}\sim\mathrm{A}(\cdot|u_{i},\Omega_{i})$ /* Analyze the current scene. */
       $h_{i}\sim\mathrm{L}(\cdot|u_{i})$ /* Retrieve required skills. */
       $x_{i}\leftarrow\textrm{GenerateCodeWithInspection}(u_{i},s_{i},h_{i})$;
       $\Omega_{i+1}\leftarrow\textrm{CompileAndRun}(x_{i},\Omega_{i})$;
end for